# An All-digital 65-nm Tsetlin Machine Image Classification Accelerator with 8.6 nJ per MNIST Frame at 60.3k Frames per Second

### v1

@@ VII Conclusion
In this work, an ASIC implementation of a ConvCoTM accelerator for inference is presented. It operates on booleanized images of 28$\times$28 pixels, and samples are classified into 10 categories. A pool of 128 clauses, i.e., pattern recognition logical expressions, is applied for the accelerator. The solution is fully programmable, and achieves a test accuracy of 97.42% on MNIST. With a clock frequency of 27.8 MHz, the accelerator’s throughput is 60.3k images per second, and with an operating voltage of 0.82 V, the energy consumed by the accelerator core per classification is 8.6 nJ. To the authors’ best knowledge this is lower than for previously reported fabricated ICs for the same classification task and for similar test accuracy.
The ASIC demonstrates the simplicity, hardware-friendliness and power-efficiency of TM solutions. We believe the TM represents an attractive ML solution for low-power edge nodes in IoT systems. Furthermore, in a TM ASIC, full on-device training capabilities can be included, enabling adaptive edge solutions.

### v2

@@ VIII Conclusion
In this work, an ASIC implementation of a ConvCoTM accelerator for inference is presented. It operates on booleanized images of 28$\times$28 pixels, and samples are classified into 10 categories. A pool of 128 clauses, i.e., pattern recognition logical expressions, is applied for the accelerator. The solution is fully programmable, and achieves a test accuracy of 97.42% on MNIST. With a clock frequency of 27.8 MHz, the accelerator’s throughput is 60.3k images per second, and with an operating voltage of 0.82 V, the energy consumed by the accelerator core per classification is 8.6 nJ. We believe the TM represents an attractive ML solution for low-power edge nodes in IoT systems for various workloads. The ASIC demonstrates the simplicity, hardware-friendliness and power-efficiency of all-digital TM solutions.